# Is this the real life? Is this just fantasy? The Misleading Success of Simulating Social Interactions With LLMs

### v2

@@ Abstract
Recent advances in large language models (LLM) have enabled richer social simulations, allowing for the study of various social phenomena with LLM-based agents. However, most work has used an omniscient perspective on these simulations (e.g., single LLM to generate all interlocutors), which is fundamentally at odds with the non-omniscient, information asymmetric interactions that humans have. To examine these differences, we develop an evaluation framework to simulate social interactions with LLMs in various settings (omniscient, non-omniscient). Our experiments show that interlocutors simulated omnisciently are much more successful at accomplishing social goals compared to non-omniscient agents, despite the latter being the more realistic setting. Furthermore, we demonstrate that learning from omniscient simulations improves the apparent naturalness of interactions but scarcely enhances goal achievement in cooperative scenarios. Our findings indicate that addressing information asymmetry remains a fundamental challenge for LLM-based agents.
@@
Figure 1: An illustration between Script simulation and Agents simulation. In the Agents simulation, two agents, each equipped with an LLM, negotiate and strategically seek information to reach a mutual agreement. Conversely, in Script simulation, a single omniscient LLM orchestrates the entire interaction based on full access to the agents’ goals. While initially appearing efficient, this interaction lacks essential human communication properties.
@@ 1 Introduction
People navigate everyday social interactions easily despite not having access to other’s mental states (i.e., information asymmetry; Weber 1978; Tomasello 1999; Oey et al. 2023). As illustrated in Figure 1, the communication between two agents that are bargaining over a price requires complex interactions for them to understand the interlocutor’s motive. With modern-day LLMs, simulating such interactions has gotten better. From building a town of AI-powered characters (Park et al., 2023) to simulating social media platforms (Park et al., 2022), and training better chatbot systems (Kim et al., 2023a; Hong et al., 2023; Chen et al., 2023a), LLMs seem to be capable to realistically simulate human social interactions.
@@
However, despite their impressive abilities, one key shortcoming has prevented realistic social simulation: a wide range of prior research has leveraged a more omniscient perspective to model and simulate social interactions (Liang et al., 2023; Li et al., 2023a; Pang et al., 2024). By generating all sides of interaction at once or making agent goals transparent to all participants, these simulations diverge from the non-omniscient human interactions that rely on social inference to achieve goals in real-world scenarios (Goodman and Frank, 2016). Studying these omniscient simulations could potentially lead to biased or wrong conclusions Das et al. (2024).
@@
To investigate the effect of this incongruity, we create a unified simulation framework by building on Sotopia (Zhou et al., 2024), a platform to simulate and evaluate social interactions with LLMs. We set up two modes for simulating human interaction in LLMs: Script mode and Agents mode. As shown in Figure 1, in the Script mode, one omniscient LLM has access to all the information and generates the entire dialogue from a third-person perspective. In the Agents mode, two LLMs assume distinct roles and engage in interaction to accomplish the task despite the presence of information asymmetry.
Comparing these two modes, we find that there are drastic disparities in each of these modes in terms of achieving social goals and naturalness. The Script mode significantly overestimates the ability of LLM-agents to achieve social goals, while LLM-based agents struggle to act in situations with information asymmetry. Additionally, the Agents mode generates interactions that sound significantly less natural, further highlighting the disparities in these simulation modes.
We then ask the question of whether LLM agents can be learned from Script simulations. Inspired by Kim et al. (2023a); Hong et al. (2023), we finetune GPT-3.5 (Ouyang et al., 2022) on a large dataset of interactions generated omnisciently. We find that through finetuning, Agents models become more natural yet barely improve in cooperative scenarios with information asymmetry. Further analysis shows that Script simulations contain information leakage in cooperative scenarios and tend to produce overly agreeable interlocutors in competitive settings.
Our findings suggest that the success of LLMs in simulating social interactions with Script mode can be misleading. While simulations generated from the third-person perspective of Script score highly in terms of goal completion rate and dialogue fluidity, the conversation strategies used by these LMs over-rely on the benefit of having direct access to the internal states of both parties. These artifacts hinder Script ability to simulate human-like interaction (Tjuatja et al., 2023; Ziems et al., 2023), and likely lead to an overestimation of the social capabilities of LLMs (Shanahan, 2023; Bender et al., 2021). Based on our findings, we provide recommendations for reporting LLM-based agent work, encouraging more careful considerations and transparency in using LLMs to simulate social interactions from both data and learning perspectives.
@@ 2 Background & Related Work
Multi-agent simulations offer an opportunity to explore complex social systems’ theories over time, serving as a generative complement to conventional empirical research methodologies (Sawyer, 2005). Previous simulation environments have been crucial for theory building and hypothesis formation across various disciplines, but they often limit agents’ communicative abilities to artificial languages and present a highly reductionist view of simulated human behavior (Gilbert, 2005; Tesfatsion and Judd, 2006; Huang et al., 2014). The recent advancements in LLMs have enabled the development of more realistic and expressive agents, which can be used to simulate social interactions in a more naturalistic manner (Park et al., 2023, 2022; Zhou et al., 2024; Li et al., 2023a). These simulations, operating in different settings and scenarios, are hard to compare. Meanwhile, some of the simulations are based on assumptions often divergent from human social interactions, which may mislead downstream applications and the public’s understanding of AI capabilities (Hendrycks et al., 2023).
@@
A common issue in training social chitchat models (i.e., chatbots) is the lack of large-scale, high-quality training data, which can be addressed by using LLMs to generate synthetic text data (Smith et al., 2020; Chen et al., 2023b). Kim et al. (2023a) first introduced SODA, a large-scale synthetic dataset for training chatbots, and showed that training chatbots with synthetic data generated by LLMs can improve their naturalness and consistency. There are also works that use LLMs to generate synthetic data (Script mode) for training chatbots in a goal-oriented setting, either using reinforcement learning (Hong et al., 2023) or using techniques to bootstrap the training data (Ulmer et al., 2024). However, these works mostly consider chitchat settings and overlook more complex scenarios involving cooperative or competitive motives. Consequently, the impact of learning from generated scripts on models’ ability to navigate complex, multi-turn interaction scenarios and accomplish social tasks remains elusive.
Modeling Human Communication
Information asymmetry is a characteristic part of human linguistic interaction (Stalnaker, 2014). It poses a challenge when we attempt to jointly achieve goals Tomasello (1999) and is exploitable in cases where one party is attempting to deceive the other (Oey et al., 2023). It also plays a large part in the human ability to achieve social goals in dialogue through strategic information omission and indirectness Pinker et al. (2008); Yoon et al. (2020); Radkani et al. (2022); Bridgers et al. (2023); Achimova et al. (2023); Carcassi and Franke (2023). Despite the asymmetry, human communication is remarkably efficient (Kemp and Regier, 2012; Zaslavsky et al., 2018; Gibson et al., 2019), and humans show a remarkable ability to track and modulate their behavior based on information added to the common ground during dialogue (Clark, 1996; Stalnaker, 2014).
@@ 3 Script vs Agents Simulation
To investigate whether the success of the omniscient Script mode reflects how LLMs would behave in the realistic human communication setting, we set up a unified framework to generate synthetic text data for different simulation settings and compare the performance of LLMs in these settings. In this section, we first introduce the general framework of agent-based simulation and Script simulation, and then we simulate social interactions across these settings to answer the following research questions (RQ): RQ1: Do the Script simulations reflect how LLMs achieve social goals in the agent-based simulation? RQ2: Do the Script simulations reflect how LLMs communicate in the agent-based simulation?
@@ 3.1 The Unified Framework for Simulation
We explore three simulation modes in our experiments. For the Script mode, one LLM has access to all the information of the characters, relationships, and social scenarios, and generates the entire social interactions at one turn from an omniscient perspective with a third-person point of view. For the Agents mode, each LLM is assigned a character and has access only to the information of the corresponding character, relationship, and social scenario. The LLMs interact with each other to complete the social task from a first-person point of view in a turn-by-turn manner. To study the effects of information asymmetry, we add one ablation setting where information asymmetry is removed from the Agents simulation by giving each agent access to other characters’ information (e.g., social goals). We refer to this setting as Mindreaders mode. [^2] Please refer to the Appendix B to see the full prompts we design for each mode.
@@
As human social behaviors are primarily driven by their social goals (Tomasello, 2021; Weber, 1978), we consider the ability to complete the social goals as one of the major indicators of the success of social interactions. Following Sotopia, we use the goal completion score (ranging from 0 to 10, higher scores indicate the agents achieve their social goals better) as the main metric to evaluate the success of the social interactions across different modes. [^3] We also evaluate using other Sotopia dimension of the social interactions (e.g., knowledge gain), and we do not observe consistent trends across different settings. Please refer to the Appendix D for more details.
@@ 3.3 RQ1: Script mode overestimates LLMs’ ability to achieve social goals
Figure 2 shows the average goal completion rate of different models in different simulation settings. We find that the Script and Mindreaders simulations achieve a significantly higher goal completion rate than the Agents simulations. This suggests that information asymmetry, which is only present in Agents mode, hinders agents’ ability to achieve social goals, and Script mode vastly overestimates that ability.
@@ 3.4 RQ2: Script mode overstates LLMs’ capability of natural interactions
Figure 3: Examples of the generated interactions from different simulation settings. All the examples are generated by GPT-3.5. We observe: (1) Script simulations contain more non-verbal communication in the simulation; (2) agent-based simulations tend to generate more repetitive utterances.
@@
Overall, our findings show that drastic disparities exist between Script and Agents simulations. Script mode overestimates LLMs’ ability to interact in realistic scenarios with information asymmetry (i.e., the Agents mode).
@@ 4.4 RQ4: Script simulations can be biased
For the cooperative scenarios, we measure the relative position of the mutual friend’s name mentioned in the conversation as a proxy for the information leakage. A value of 0 indicates the name was mentioned at the start of the conversation, while a value of 1 indicates it was mentioned at the end. Script mode results show an average first-mention location of 0.13, contrasting with Agents mode, which has an average of 0.39. This suggests that in Script mode, the mutual friend’s name is ‘guessed’ almost immediately. The complete distribution is in Figure 12 in the Appendix. This demonstrates a bias of Script mode exploiting its knowledge from the omniscient perspective about the conversational participants. We find that this strategy generalizes poorly to the setting where models do not have ground truth access to their interlocutor’s knowledge and goals (as shown in Figure 6). This aligns with recent findings that LLM abuses its omniscient perspective in information-asymmetric contexts Kim et al. (2023b).
@@ 5 Conclusion & Discussion
We scrutinize recent advances in social simulation by evaluating current approaches’ ability to generalize to settings that are closer to human interaction. Focusing on cooperation and competition given information-asymmetric settings, we evaluate three modes of deploying LLMs based on past approaches in the literature. We find that while Script mode, the widely-used method for social simulation, achieves natural simulations but shows a bias toward exploiting white box access to the participants early in the interaction. Furthermore, we find that finetuning models on these generations improve selectively on a measure of goal completion from Sotopia, but it also imbues the implausible strategies from the ‘omniscient’ Script simulations into the student models, resulting in further bias. Below, we discuss the implications of our findings around the limitations of omniscient perspective (§5.1), and provide concrete recommendations for reporting and evaluating the results of LLM simulations (§5.2), and conclude by proposing potential approaches for improving LLM based conversational agent simulations (5.3).
@@ 5.2 Recommendations for Reporting
5.3 Towards More Realistic Social Simulations
@@
While it is plausible that future models will improve on one or both of these axes with increased scale, current interaction simulation could benefit from structuring generations to provide models with more human-like access to their interlocutor’s mental state. One mechanism for this could be meticulous data curation to prevent the models from ‘hacking’ them via shallow heuristics (Hong et al., 2023; Ulmer et al., 2024). Another promising avenue is prompting language models to cooperatively build an explicit text-based log of the conversational common ground, described in Stalnaker (2014) as the “evolving body of background information that is presumed to be shared by the participants in a conversation.”

### v3

@@ Abstract
Recent advances in large language models (LLM) have enabled richer social simulations, allowing for the study of various social phenomena. However, most recent work has used a more omniscient perspective on these simulations (e.g., single LLM to generate all interlocutors), which is fundamentally at odds with the non-omniscient, information asymmetric interactions that involve humans and AI agents in the real world. To examine these differences, we develop an evaluation framework to simulate social interactions with LLMs in various settings (omniscient, non-omniscient). Our experiments show that LLMs perform better in unrealistic, omniscient simulation settings but struggle in ones that more accurately reflect real-world conditions with information asymmetry. Our findings indicate that addressing information asymmetry remains a fundamental challenge for LLM-based agents.
@@
Figure 1: An illustration between Script mode simulation and Agents mode simulation. In the Agents mode, two agents, each equipped with an LLM, negotiate and strategically seek information to reach a mutual agreement. Conversely, in Script mode, a single omniscient LLM orchestrates the entire interaction based on full access to the agents’ goals. These two modes end up on opposite sides of the spectrum in terms of information asymmetry from various perspectives (e.g., roles, social goals, secrets, etc.).
@@ 1 Introduction
LLM-based social simulations have become great tools for modeling human behavior in various contexts (Park et al., 2023; Sreedhar and Chilton, 2024), understanding and measuring LLMs’ social skills through certain tasks (Zhou et al., 2024; Bianchi et al., 2024), improving and aligning chatbot systems by providing training data (Kim et al., 2023a; Hong et al., 2023; Chen et al., 2023a; Pang et al., 2024). The success in replicating human-like social interactions across diverse domains offers a promising picture of the realistic social capabilities of LLMs.
@@
However, the role of information asymmetry in these simulations, i.e., the degree to which interlocutors in interactions have access to each other’s internal private mental states and goals, has been largely overlooked (Weber, 1978; Tomasello, 1999; Oey et al., 2023) [^1] We extend the scope of the traditional definition of information asymmetry to encompass broader social aspects.. Instead of using the more realistic simulation setting that mirrors human daily social interactions with information asymmetry (e.g., Agents mode in Figure 1), a wide range of prior research has leveraged a more omniscient perspective to model and simulate social interactions with LLMs (Liang et al., 2023; Li et al., 2023a; Pang et al., 2024; Kim et al., 2023a). By generating all sides of interaction at once or making agent social goals or tasks transparent to all participants, these simulations diverge from the non-omniscient human interactions that rely on social inference to achieve goals in real-world scenarios (Goodman and Frank, 2016). Studying these omniscient simulations could lead to biased or wrong conclusions about LLMs’ social capabilities Das et al. (2024).
@@
To investigate the effect of this incongruity, we create a unified simulation framework with two distinct modes for simulating human interaction with LLMs: Script mode and Agents mode. As shown in Figure 1, in the Script mode, one omniscient LLM has access to all the information and generates the entire dialogue from a third-person perspective (e.g., Kim et al. 2023a; Chen et al. 2023b). In the Agents mode, two LLM agents assume distinct roles and engage in interaction to accomplish the task (e.g., Zhou et al. 2024). These modes represent the opposite ends of the spectrum regarding information asymmetry, while the Agents mode is the realistic interaction simulation setting that reflects the information asymmetry in human daily-life interactions.
We first compare the interactions produced in these two simulation modes, examining the extent to which the simulated characters achieve their social goals at the end of the interaction, as well as the naturalness of the interactions. We find that LLMs in the Agents mode not only struggle to generate social interactions that effectively meet the specified social goals for each role but also produce less naturally flowing social interactions, particularly in their utterances when compared to the LLMs in the Script mode. These findings indicate that LLMs still fall short of acting as agents and simulating social interaction within contexts of realistic human interaction settings.
We then ask the question of whether LLM agents can be learned from Script simulations. Inspired by Kim et al. (2023a); Hong et al. (2023), we finetune GPT-3.5 (Ouyang et al., 2022) on a large dataset of interactions generated in the Script mode. We find that finetuning on omnisciently generated social interactions provides limited improve for LLMs interacting in the Agents mode. Further data analysis reveals the biases within Script mode simulations, hindering the ability of models trained on such data to effectively generalize real-world social skills.
Based on our findings, we provide recommendations for reporting LLM-based agent work, encouraging more careful considerations and transparency in using LLMs to simulate social interactions from both data and learning perspectives.
@@ 2 Background & Related Work
Realistic, humanlike simulation settings have been crucial for social theory building and hypothesis formation across various disciplines (Gilbert, 2005; Tesfatsion and Judd, 2006; Huang et al., 2014). The recent advancements in LLMs have enabled the development of social simulations driven by human language (Park et al., 2023, 2022; Zhou et al., 2024; Li et al., 2023a). However, these LLM-based simulations often operate in settings divergent from human social interactions, which may mislead downstream applications and the public’s understanding of AI capabilities (Hendrycks et al., 2023). Furthermore, many of these works lack a consistent evaluation framework, while SOTOPIA (Zhou et al., 2024) has begun addressing this gap by offering a holistic evaluation framework for assessing social interactions generated by LLMs.
@@
A common issue in training social chitchat models (i.e., chatbots) is the lack of large-scale, high-quality training data, which can be addressed by using LLMs to generate synthetic text data (Smith et al., 2020; Chen et al., 2023c). Kim et al. (2023a) first introduced SODA, a large-scale synthetic dataset for training chatbots to produce more natural and consistent utterances. There are also works that use LLMs to generate synthetic data (Script mode) for training chatbots in a goal-oriented setting, either using reinforcement learning (Hong et al., 2023) or using techniques to bootstrap the training data (Ulmer et al., 2024). However, these works mostly consider chitchat settings and overlook more complex scenarios involving cooperative or competitive motives. Consequently, the impact of learning from generated scripts on models’ ability to navigate complex, multi-turn interaction scenarios and accomplish social tasks remains elusive.
Information Asymmetry in Communication
Information asymmetry is a characteristic part of human linguistic interaction (Stalnaker, 2014). It poses a challenge when we attempt to jointly achieve goals Tomasello (1999) and is exploitable in cases where one party is attempting to deceive the other (Oey et al., 2023). It also plays a large part in the human ability to achieve social goals in dialogue through strategic information omission and indirectness Pinker et al. (2008); Yoon et al. (2020); Radkani et al. (2022); Bridgers et al. (2023); Achimova et al. (2023); Carcassi and Franke (2023). In LLM-driven social simulations, information asymmetry is examined through the variability in prompts provided to each generation iteration. This incorporates a range of factors including assigned roles (e.g., assistant or user), specific output restrictions (e.g., "only ask questions"), character backgrounds (e.g., "you are a doctor"), and particular social objectives (e.g., "your goal is to borrow \$2000"). The varied elements unique to each agent help simulate the complexities and nuances of real-life social interactions within the framework of the simulation.
@@ 3 Script vs Agents Simulation
To investigate whether the success of the omniscient Script mode reflects how LLMs would behave in the realistic human communication setting, we set up a unified framework to generate synthetic text data for different simulation settings and compare the performance of LLMs in these settings. In this section, we first introduce the general framework of agent-based simulation and Script simulation, and then we simulate social interactions across these settings to answer the following research questions (RQ): RQ1: Do the Script simulations reflect how LLMs achieve social goals in the realistic soical interaction settings? RQ2: Do the Script simulations reflect how LLMs communicate in the realistic soical interaction settings?
@@ 3.1 The Unified Framework for Simulation
We explore three simulation modes in our experiments. For the Script mode, one LLM has access to all the information of the characters, relationships, and social scenarios, and generates the entire social interactions at one turn from an omniscient perspective with a third-person point of view. For the Agents mode, each LLM is assigned a character and has access only to the information of the corresponding character, relationship, and social scenario. The LLMs interact with each other to complete the social task from a first-person point of view in a turn-by-turn manner. Note that unlike other previous works that only have one or two sources of information asymmetry (e.g., occupation; Pang et al. 2024), our Agents mode simulation can have a diverse array of asymmetrical factors, including gender, age, occupation, personality, secretive information, and social goals. To further study the effects of information asymmetry, we add one ablation setting where each agent has access to other characters’ information (e.g., social goals and secretive information). We refer to this setting as Mindreaders mode. [^3] Please refer to the Appendix B to see the full prompts we design for each mode.
@@
As human social behaviors are primarily driven by their social goals (Tomasello, 2021; Weber, 1978), we consider the ability to complete the social goals as one of the major indicators of the success of social interactions. Following Sotopia, we use the goal completion score (ranging from 0 to 10, higher scores indicate the agents achieve their social goals better) as the main metric to evaluate the success of the social interactions across different modes. [^4] We also evaluate using other Sotopia dimension of the social interactions (e.g., knowledge gain), and we do not observe consistent trends across different settings. Please refer to the Appendix D for more details. Note that the goal completion score is a proxy for the success of the social interactions, and we use model-based evaluation to obtain the esitmation of the goal completion score following Zhou et al. (2024).
@@ 3.3 RQ1: Script mode overestimates LLMs’ ability to achieve social goals
Figure 2 shows the average goal completion rate of different models in different simulation settings. We find that the Script and Mindreaders simulations achieve a significantly higher goal completion rate than the Agents simulations. This suggests that information asymmetry hinders agents’ ability to achieve social goals, and Script mode vastly overestimates LLMs’ ability to achieve social goals in realistic, humanlike social interaction settings.
@@ 3.4 RQ2: Script mode overstates LLMs’ capability of natural interactions
Figure 3: Illustrative examples of the generated interactions from different simulation settings. All the examples are generated by GPT-3.5. Note that our actual prompts are more complex than the content in the green box (see Appendix B). We observe: (1) Script simulations contain more non-verbal communication in the simulation; (2) agent-based simulations tend to generate more repetitive utterances.
@@
Overall, our findings show that drastic disparities exist between Script and Agents simulations. Script mode overestimates LLMs’ ability to interact in realistic settings with information asymmetry (i.e., the Agents mode).
@@ 4.4 RQ4: Script simulations can be biased
For the cooperative scenarios, we measure the relative position of the mutual friend’s name mentioned in the conversation as a proxy for the information leakage. A value of 0 indicates the name was mentioned at the start of the conversation, while a value of 1 indicates it was mentioned at the end. Script mode results show an average first-mention location of 0.13, contrasting with Agents mode, which has an average of 0.39. This suggests that in Script mode, the mutual friend’s name is ‘guessed’ almost immediately. The complete distribution is in Figure 12 in the Appendix. This demonstrates a bias of Script mode exploiting its knowledge from the omniscient perspective about the conversational participants. We find that this strategy generalizes poorly to the setting where models do not have ground truth access to their interlocutor’s knowledge and goals (as shown in Figure 6).
@@ 5 Conclusion & Discussion
We scrutinize recent advances in social simulation by evaluating current approaches’ ability to generalize to settings that are closer to human interaction. Focusing on cooperation and competition given information-asymmetric settings, we evaluate three modes of deploying LLMs based on past approaches in the literature. We find that LLMs continue to face challenges when operating in more realistic Agents mode. Meanwhile, the simulations generated from the Script mode show biases toward exploiting white box access to the participants early in the interaction. Furthermore, we find that finetuning models on these generations improve selectively on a measure of goal completion from Sotopia, but it also imbues the implausible strategies from the ‘omniscient’ Script simulations into the student models, resulting in further bias.
@@ 5.2 Recommendations for Reporting
5.3 Towards Better Simulations in More Realistic Settings
@@
While it is plausible that future models will improve on one or both of these axes with increased scale, current interaction simulation could benefit from structuring generations to provide models with more human-like access to their interlocutor’s mental state. One possible solution is meticulous data curation to thwart models from exploiting shallow heuristics (Hong et al., 2023; Ulmer et al., 2024). Another approach involves prompting language models to collaboratively construct an explicit text-based log of the shared conversational context, as described by Stalnaker (2014).
@@ 6 Limitations and Ethical Considerations
Promt Design
Our work is built on the prompt framework in (Zhou et al., 2024) to simulate social interactions. The prompts contain multiple structured fields, such as the role of each agent, the goal of the interaction, and the constraints on the interaction. We acknowledge that the prompt design may not fully capture the complexity of human social interactions, and switching to different simulation frameworks with different prompt designs may lead to variations in the results. However, the main goal of this work is to reveal the challenges of realistically simulating social interactions with LLMs due to information asymmetry. And such challenges are likely to persist across different prompt designs. Future work should explore how different prompt designs affect the performance of LLMs in social simulations.